# Revisiting Pre-trained Language Models for Vulnerability Detection

### v1

@@ IV-F RQ6: Are PLMs capable of detecting vulnerable code after semantic-preserving transformation?
[…]
V Threats to Validity
Internal Validity. To minimize data leakage, we limit our dataset to commits dated after the cutoff date of the evaluated LLMs. Although some undetectable forms of leakage may persist, addressing them is beyond the scope of this study. We excluded Chain-of-Thought (CoT) prompting due to budget constraints. However, prior studies have shown that the effectiveness of CoT varies across specific scenarios [20, 32, 19]. Therefore, this exclusion does not undermine the significant performance gap between fine-tuning and prompt engineering observed in our findings. In practice, the decision between fine-tuning and prompting is still a trade-off between cost and performance, determined by user requirements.
External Validity. This work focuses on evaluating PLMs for function-level VD, as mainstream PLMs with limited context window sizes are not yet sufficient for addressing these more complex cross-function VD. We exclude latest reasoning models (e.g., OpenAI o3) due to the limited availability of new vulnerability data for evaluation and the potential risk of data leakage. However, our extensible data collection framework supports future evaluations using up-to-date CVEs patches.
VI Conclusion
This work presents an extensive evaluation of the vulnerability detection capabilities of representative pre-trained language models, covering Code SLMs, Code LLMs, and general LLMs. Among our findings, we identify that: PLMs incorporating pre-training tasks designed to capture the syntactic and semantic patterns of code outperform both general-purpose PLMs and those solely pre-trained or fine-tuned on large code corpora; With code structure-aware pre-training tasks, PLMs are robust to code abstraction; However, they remain sensitive to code normalization and transformation, and face challenges in handling vulnerabilities with complex dependencies or those exceeding the context window size.
Appendix A Prompt Setting/Type Impact on LLM Performance
Figure 6: LLM performance by prompt settings and types
@@ Appendix A Prompt Setting/Type Impact on LLM Performance
In addition to the best performance of LLMs across all prompt settings and types reported in our prompting-based evaluations (Section IV-B), this section provides a detailed analysis of how different prompt configurations influence model performance. As shown in Fig. 6, with respect to prompt types, GPT-4o Mini achieves its highest performance when API calls are incorporated into the prompts. The effectiveness of adding supplementary structural information to prompts depends heavily on the model’s inherent code understanding capabilities. For instance, models like DeepSeek-Coder typically benefit from the inclusion of API call and data flow, as such additions help reinforce code logic and clarify dependencies. This is likely because DeepSeek-Coder already possesses a solid understanding of raw code. On the other hand, Code LLMs such as Code Llama and StarChat-$\beta$, which perform at or below random guessing on raw code, often struggle when confronted with these added logic chains, making code comprehension even more challenging.
@@
Regarding prompt settings, few-shot in-context learning enhances the performance of LLMs (e.g., GPT-4o Mini) by leveraging extended context windows and reference examples. This approach serves as implicit guidance, encouraging models to follow instructions more effectively [75]. However, for Code LLMs such as DeepSeek-Coder and WizardCoder, which already show strong performance in zero-shot scenarios, introducing a limited number of few-shot examples may disrupt decision-making [76]. The complexity and variability of vulnerabilities can make a small set of few-shot examples insufficient to convey the necessary information. Even worse, it may lead the model to overfit on simplistic patterns from provided examples, rather than learning generalizable features.
[…]

### v3

@@ 4.6. RQ6: Are PLMs capable of detecting vulnerable code after semantic-preserving transformation?
4.8. RQ8: How do PLMs perform under different prompt types and settings?
Figure 10. PLM performance by prompt settings and types:
@@ 4.8. RQ8: How do PLMs perform under different prompt types and settings?
In addition to the best performance of LLMs across all prompt settings and types reported in our prompting-based evaluations (Section 4.2), this section provides a detailed analysis of how different prompt configurations influence model performance. As shown in Fig. 10, with respect to prompt types, GPT-4o Mini achieves its highest performance when API calls and few-shot examples are incorporated into the prompts. The effectiveness of adding supplementary structural information to prompts depends heavily on the model’s inherent code understanding capabilities. For instance, models like DeepSeek-Coder typically benefit from the inclusion of API call and data flow, as such additions help reinforce code logic and clarify dependencies. This is likely because DeepSeek-Coder already possesses a solid understanding of raw code. On the other hand, Code LLMs such as CodeLlama and StarChat-$\beta$, which perform at or below random guessing on raw code, often struggle when confronted with these added logic chains, making code comprehension even more challenging.
@@
Regarding prompt settings, few-shot in-context learning enhances the performance of LLMs (e.g., GPT-4o Mini) by leveraging extended context windows and reference examples. This approach serves as implicit guidance, encouraging models to follow instructions more effectively (ICT). However, for Code LLMs such as DeepSeek-Coder and WizardCoder, which already show strong performance in zero-shot scenarios, introducing a limited number of few-shot examples may disrupt decision-making (ACL). The complexity and variability of vulnerabilities can make a small set of few-shot examples insufficient to convey the necessary information. Even worse, it may lead the model to overfit on simplistic patterns from provided examples, rather than learning generalizable features.
Answer-8: The effectiveness of different prompt settings and types depends on a model’s inherent code understanding. Models like GPT-4o Mini benefit from added structural information when they already exhibit a fundamental understanding of raw code, while more complex prompts may hinder weaker models. Given the complex and diverse vulnerability patterns that are difficult to capture with limited few-shot examples, in-context learning may mislead the model to overfit superficial patterns.
5. Discussion
This paper differentiates itself from previous work by revisiting the true capabilities of PLMs for VD, addressing the shortcomings of existing VD evaluations. We argue that the conclusions drawn from previous, biased evaluations are meaningless, as they lead to a significant gap between reported performance and practical effectiveness. This is a key point emphasized throughout our work.
For dataset selection, to minimize data leakage, we limit our dataset to commits dated after the knowledge cutoff dates of all evaluated LLMs. While some undetectable forms of leakage may still exist, addressing them is beyond the scope of this study.
For model selection, we acknowledge that some open-source reasoning models (e.g., Qwen3 (qwen3)) have shown strong performance in the software engineering tasks. However, their very recent knowledge cutoff date conflicts with the scarcity of latest vulnerabilities, making it difficult to collect enough recent data to ensure the accuracy and comprehensiveness of evaluation results. We also clarify that while resource and budget constraints prevent us from evaluating all SOTA LLMs, our evaluation framework can help future research evaluate stronger LLMs on more recent VD data.
Although our findings conclude that Code SLMs (e.g., PDBERT), which are trained on specialized code pretraining tasks, generally outperform LLMs trained on general language modeling, their applicability is constrained to function-level VD due to the limited context window size. In contrast, LLMs provide a larger context window, but their ability to perform longer function-level or cross-procedure VD still needs significant improvement.
6. Conclusion
This work presented an extensive evaluation of the vulnerability detection capabilities of representative PLMs. The findings revealed that PLMs incorporating pre-training tasks designed to capture the syntactic and semantic patterns of code outperform those solely pre-trained on general language modeling tasks. However, they remain sensitive to code normalization and transformation, and face challenges in handling vulnerabilities with complex dependencies or those exceeding the context window size.
Appendix A Ethical Considerations
Our work evaluates existing pre-trained language models on the vulnerability detection task. All experiments are conducted using publicly available data, models, and source code. Since no new vulnerabilities are generated or identified in this work, no ethical issues will arise.
Appendix B Models
B.1. Small Language Models for Code (Code SLMs)
CodeBERT (codebert) is a bimodal pre-trained encoder-based Transformer model that utilizes a multi-layer bidirectional Transformer-based architecture to learn general-purpose representations for programming language (PL) and natural language, achieving good performance in the natural language code search and code documentation generation through incorporating a replaced token detection objective.
GraphCodeBERT (graphcodebert) is a pre-trained encoder-based Transformer model that incorporates inherent semantic structure of code using data flow rather than traditional syntactic structures like ASTs to enhance code understanding. By employing a multi-layer bidirectional Transformer-based architecture and introducing two structure-aware pre-training tasks (i.e., edge prediction and node alignment), it achieves satisfactory performance on tasks such as code search, clone detection, code translation, and code refinement.
UniXCoder (UniXcoder) is a unified cross-modal pre-trained model designed for PL, which optimizes code understanding and generation tasks by utilizing mask attention matrices and prefix adapters, as well as integrating cross-modal contents such as AST and code comments to achieve code fragment representation learning. It demonstrates advanced performance across five code-related tasks and a new zero-shot code-to-code search task, highlighting the benefits of leveraging both comments and AST structures in its architecture.
PDBERT (pdbert) is a pre-trained encoder-based Transformer model designed to enhance dependency analysis in software by introducing two novel pre-training objectives: statement-level control dependency prediction and token-level data dependency prediction. PDBERT significantly improves understanding of vulnerable code and achieves strong performance on vulnerability detection, classification, and assessment tasks as well as program dependence analysis.
CodeT5 (codet5) is a unified pre-trained encoder-decoder Transformer model designed for effectively leveraging code semantics from developer-assigned identifiers. It introduces an identifier-aware pre-training task to differentiate and recover identifiers and utilizes user-written code comments in a bimodal dual generation task for better alignment between nature language and PL, resulting in significant performance improvements over previous methods in code understanding tasks (e.g., code defect detection and clone detection) and various generation tasks.
B.2. Large Language Models for Code (Code LLMs)
CodeLlama (codellama) is initialized from LLaMA 2 and featured by a multitask objective combining autoregressive and causal infilling prediction, enabling various tasks such as code completion and docstring generation. It captures long input contexts by adjusting the parameters of RoPE positional embeddings. Further fine-tuned on a blend of proprietary instruction data and a machine-generated self-instruct dataset, the helpful and safe CodeLlama-Instruct models better follow human instructions and achieve excellent performance among open models on several code benchmarks, including HumanEval and MBPP.
DeepSeek-Coder (deepseekcoder) is initialized from the DeepSeek LLM. By pre-training with a fill-in-the-middle approach on a combination of project-level source code and natural language corpora in English and Chinese, DeepSeek-Coder models enhance their versatility and applicability across various coding scenarios. Extending the input sequence length by reconfiguring RoPE parameters further strengthens the models’ reliable ability to handle long-context tasks, such as cross-file code completion. Incorporating high-quality, helpful, and impartial human instruction data, DeepSeek-Coder-Instruct significantly narrows the performance gap between OpenAI’s GPT-4 and open-source models on a wide range of code-related benchmarks.
StarChat-$\beta$ (starchat) is initialized from StarCoderBase, which was pre-trained on the Stack dataset covering over 80 programming languages. Fine-tuning StarCoderBase on 35B Python tokens enables StarCoder to perform well in Python and other languages. To further enhance its capabilities as a coding assistant, StarChat-$\beta$ is fine-tuned from StarCoderPlus, which was trained on an uncensored variant of the OpenAssistant-Guanaco dataset, making the model more helpful for coding tasks.
WizardCoder (wizardcoder) is initialized from the StarCoder series. It applies the Evol-Instruct method, proposed by WizardLM, to construct the instruction-following training set for fine-tuning. Comprehensive experiments on four code generation benchmarks (i.e., HumanEval, HumanEval+, MBPP, and DS1000) reveal the model’s exceptional capabilities compared to other open-source code LLMs.